Circular No. 6372

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) D. Schleicher, Lowell Observatory, communicates: “Narrowband pho-

tometry of comet C/1996 B2 was obtained by D. Osip and S. Lederer (University of Florida) using the 0.79-m telescope at Lowell Observatory on Mar. 23.2–23.5 UT (r = 1.08 AU), Mar. 24.2–24.5 (1.06 AU), and Mar. 25.1–25.2 (1.05 AU). Mean Haser-model production rates for the three dates, respectively, were: log Q(OH) = 29.17, 29.18, and 28.93; log Q(CN) = 26.69, 26.67, and 26.60; log Q(C_{2}) = 26.86, 26.84, and 26.78. The dust-production rates were log (Afρ) = 3.91, 3.82, and 3.75. Additional photometry was obtained by R. Millis, D. Schleicher, and L. Wasserman using the Hall 1.07-m telescope at Lowell Observatory on Mar. 29.1–29.3 (0.96 AU) and 30.1–30.2 (0.94 AU). Production rates for these dates were: log Q(OH) = 28.88 and 28.89; log Q(CN) = 26.56 and 26.58; log Q(C_{2}) = 26.73 and 26.77; and log (Afρ) = 3.80 and 3.85. Combining these results with those previously obtained (cf. IAUC 6344) indicates that production was systematically high on Mar. 23 and 24, particularily for OH, and returned to near-baseline slopes on the Mar. 25, 29, and 30. The high values on Mar. 23 and 24 may be related to the ‘luminous knots’ reported by Lecacheux et al. (IAUC 6354) on Mar. 23.9–24.2. Well-defined periodic brightness variations were observed on all five nights for all species, with a typical amplitude in the log of ∼ 0.1 each night. Single-peaked lightcurves with periods of ∼ 6.25 or 8.55 hr are viable, while double-peaked lightcurves with periods of 9.80, 12.45, 13.0, or 17.2 hr are also acceptable. Periods shorter than ∼ 5 hr are ruled out. The single-peaked periods are slightly longer than the periods proposed by Lecacheux et al. (cf. IAUC 6354). Narrowband CCD imaging by D. Schleicher, R. Millis, and L. Wasserman using the Hall telescope on Mar. 23, 24, and 25 shows distinct morphological structure varying on a timescale of well less than 1 hr. Preliminary analysis indicates the morphology repeats with a period of between 6.0 and 6.4 hr within an individual night, while the structure does not repeat for any other period less than 10 hr. The imaging results, combined with the lightcurve results, strongly imply that the comet has a rotation period of ∼ 6.25 hr and that a single active region controls the photometric variations.”

Total visual magnitude estimates: Apr. 2.81 UT, 2.2 (H. Luethen, Hamburg, Germany, 10×50 binoculars); 3.99, 2.1 (R. J. Bouma, Groningen, Netherlands, naked eye; tail 18^{o} long; during total lunar eclipse); 4.00, 2.3 (J. D. Shanklin, Brinkley, England, naked eye; tail 7^{o} long in p.a. 44^{o}).

1996 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green